A solution is made by mixing 223.0 mL of ethanol initially at15.3 ∘C with 223.0 mL of water initially at 21.1 ∘C What is the final temperature of the solution assuming that no heat is lost? The density of ethanol is 0.789 g/mL and the density of water is 1.00 g/mL. The specific heat of ethanol is 2.46 J/g⋅°C and the specific heat of water is 4.184 J/g⋅°C.

Tf=

∘C
Answer #1

mass of ethanol = 223.0 ml * 0.789 g / ml = 175.95 g

and

mass of water = 223.0 g

let  final temperature of the solution = T oC

heat gain by ethanol = heat lost by water

or

m1 * s1 * dT1 = m2 * s2 * dT2

or

175.95 *2.46 * (T - 15.3) = 223.0 * 4.184 * (21.1 - T)

or

(T - 15.3) = 2.1557 * (21.1 - T)

or

T = 19.3 oC

thus

the final temperature is 19.3 oC